Status Report

NASA Hubble Space Telescope Daily Report #4861

By SpaceRef Editor
June 9, 2009
Filed under , ,

HUBBLE SPACE TELESCOPE – Continuing to collect World Class Science


PERIOD COVERED: 5am June 5 – 5am June 8, 2009 (DOY 156/0900z-159/0900z)


COS/NUV 11467

NUV Detector Internal Functionality and Operation

The functionality and operation of the COS NUV detector will be tested using internal sources. Several science exposures will be taken in TIME-TAG using both the DEUTERIUM and Pt/Ne lamps. This is SMOV Activity COS-06

COS05/NUV 11466

NUV Detector Dark

Measure the NUV detector dark rate by taking long science exposures with no light on the detector. The detector dark rate and spatial distribution of counts will be compared to pre-launch data in order to verify the nominal operation of the detector, and for use in the CalCOS calibration pipeline. Variations of count rate as a function of orbital position will be analyzed to find dependence of dark rate on proximity to the SAA.

FGS/S/C 11458

Optimizing FGS2R2 Performance with the AMA

This proposal satisfies SMOV4 AD OTA/FGSS-054. S-curves with PUPIL and F583W will be obtained at selected locations within the FGS2R2 FOV during SMOV4. These data will be will be used to adjust the AMA so as to align the HST pupil onto the interferometer’s Koesters prisms. The final AMA position will be set to produce an acceptable set of S-curves across the FGS2R2 FOV. The AMA will be adjusted using CCL/instructions from the ground. It is intended that this proposal executes during the SMOV4 BEA and will require several (on the order of 4) iterations to achieve optimal interferometric performance.

STIS07/CCD 11404

CCD Dark and Bias Monitor for SMOV4/Cycle 17 Activity STIS-07

Monitor the darks for the STIS CCD. Monitor the bias in the 1×1, 1×2, 2×1, and 2×2 bin settings at gain=1, and 1×1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns.

STIS13/CCD 11388

STIS-13 CCD Spectroscopic Image Quality and ACQ Tests

Check image quality along the 52×0.1 slit in combination with the G230LB grating. The target star will be positioned at several positions along the slit including 52X0.1 (nominal center position), 52X0.1E1 (CTE position) and 51X0.1D1 (MAMA acquisitions near low background for FUV MAMA). Positioning of the target will be done utilizing PATTERN and offsets of 0.1 arcseconds between five positions. Sufficient targets will be selected, and contingency visits prepared, to allow this activity to occur any time during the BEA period.

ACS/SBC 11372

SBC UV Contamination Check

ACS/SBC throughput in the far-UV will be monitored during the bright earth avoidance period through use of two non-BEA visits, each of one orbit separated by one week allowing observations of NGC 6681. A waiver for two non-BEA orbits during BEA was previously granted. Exposures will be taken in PR110L, F115LP, F125LP, F140LP, F150LP and F165LP as in recent UV Monitor program observations. These results will be one of the activities determining if the BEA can be ended. A third visit just after BEA ends will repeat the above monitoring and include one orbit with F122M. ON HOLD visits 4 and 5 would substitute for 1 and 2 if SM4 is delayed.

COS23/S/C/FUV 11356

COS FUV Initial On-Orbit Turn-On

This proposal specifies the procedure for SMOV initial HV turn-on and ramp-up of the COS FUV detector. (The FUV will have been commanded to its Operate state to support execution of proposal 11353.) The procedure is detailed in the Observing Description, but in summary, the following is done: The initial transition from FUV Operate to HVLow is broken into two parts, with a gap of 4 hours between turning on the HV and ramping to the HVLow (SAA) voltage. This will be followed by 5 cycles of HV ramp-up and return to HVLow. Cycles will ramp up to successively higher (magnitude) voltage, with the fifth cycle going to the nominal operating values. There will be a gap of at least 4 hours between cycles. All HV ramp-up will be done at 10 sec per HV “step”. The step rate and cycle voltage values (for Segments A and B) must be patched in FSW in each cycle prior to the HV ramp commanding. Memory monitors will be set on the patched memory locations. Immediately after any HV commanding, and 4 hours after ramp-up commanding, the DCE memory will be dumped. Immediately after HV ramp-up commanding higher than HVLow, short DARK & WAVE exposures will be obtained. Visits 01 and 02, and all the subsequent even numbered visits (the ones 4 hours after HV ramp-ups), end with NSSC-1 COS event flag 3 being set. If the flag remains set, subsequent FUV commanding will be skipped. Thus, Operations Requests must be in place to clear the flag prior to those subsequent visits. Real-time monitoring of the telemetry will be used to guide the decisions whether or not to clear the flag. The final visit (13) provides a scheduled final opportunity to clear flag 3, and if the flag is cleared, initiates nominal FUV HV commanding and requests a DARK exposure.

STIS17/MA/S/C 11350

FUV MAMA HV Recovery

This proposal addresses the concerns over Cesium migration from the FUV-MAMA photocathode into the pores of the microchannel plate and lays out the procedures necessary to permit a safe and controlled recovery of the FUV-MAMA detector to science. This procedure should only be used during the SMOV4 high voltage recommissioning phase for the FUV-MAMA. After that, the normal ramping procedure should be used. This proposal is loosely based on proposal 10036, STIS MAMA Anomalous Recovery, with the differences summarized below: (1) additional plateaus during voltage commanding, (2) longer dwells after each 50 V step, (3) intermediate ramp re-structured into 2 parts, (4) change in yellow and red limits and the MCP Hi limit, setting new values for the next plateau, and (5) adding a fold test at 300V below the nominal MCP voltage. The recovery consists of four separate and unique procedures (visits) and they must be completed successfully and in order. They are: (1) a signal processing electronics check, (2) 1st high voltage ramp-up to an intermediate MCP voltage of -1500 V with limits modifications and voltage plateaus, (3) 2nd high voltage ramp-up to an intermediate MCP voltage of -1950 V (300 V below the nominal MCP voltage) with limits modifications and voltage plateaus followed by a fold distribution test, and (4) the final high voltage ramp-up to the full operating voltage, again with limits modifications and voltage plateaus, followed by a fold distribution test. During the 1st high voltage ramp-up, diagnostics are performed followed by a dark exposure. During the 2nd and 3rd high voltage ramp-ups, diagnostics are performed followed by a dark, flat field ACCUMs, and a fold analysis test. All required parameters are defined in 07-248 FUV and NUV SMOV Ramp-up Parameters Excel workbook. Supports Activity STIS-17


Significant Spacecraft Anomalies: (The following are preliminary reports of potential non-nominal performance that will be investigated.)


11867 – At 157/13:04:22z GSAcq (1,3,3) scheduled from 13:00:29z – 13:07:30z resulted in Fine Lock Back-up (1,0,1).

Observations affected: None

11869 – GSAcq(1,3,3) scheduled at 156/23:23:33z – 23:31:04 resulted in fine lock backup (3,0,3) using FGS-3 due to Scan Step Limit Exceeded on FGS-1.

Observations possibly affected: COS 8 – 12 Proposal ID#11467 and Astrometry Proposal ID#11458


18474-0 – 2nd AMA Move for FGS2R2 @ 156/0951z

18490-0 – Clear COS Event Flag 3 after FUV initial turn on Part 1 completes @ 156/1330z

18493-0 – Uplink OBSINT Gyro Bias Update @ 156/1542z

18496-1 – Perform COS Time code update @ 157/1413

18495-0 – Clear COS Event Flag 3 after FUV initial turn on Part 2 completes @ 157/1426z

18468-0 – Clear STIS Event Flag 2 (MAMA1 Checkout Part 4) @ 157/1448z

18497-0 – 3rd AMA Move for FGS2R2 @ 158/2239z



FGS GSAcq 25 25
FGS REAcq 1 1
OBAD with Maneuver 18 18


Flash Reports:

-The 2nd FGS2R2 AMA Move was executed at 156/09:39-09:52z.

-The COS FUV HV was ramped to HVlow (-4069 volts) on both segments A and B at 156/16:55z

-At 157/20:54z the COS FUV HV was partially ramped to HVNom (-4916/-4869 volts) on both segments A and B respectively.

-The 3rd FGS2R2 AMA Move was executed at 158/22:30-22:39z

SpaceRef staff editor.